40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4087     Entry time: Wed Dec 22 15:30:47 2010
Author: Osamu 
Type: Update 
Category: ASC 
Subject: ETMY oplev fixed 

According to c1scy.mdl, OL signals should be connected to adc_0_24 to adc_0_27 but they were connected to adc_0_16 to adc_0_19 which are assigned to QPD signals.

Actually cable connections were messed up. One ribbon cable was connected from QPD driver and ADC ports assigned for OL, and another ribbon cable was connected from the board combining the signals of oplev and QPD to ADC port assigned for QPD.

Now ETMY oplev is working well and aligned to center.

ELOG V3.1.3-